# Generalized Correlation Regression for Disentangling Dependence in Clustered Data

### v1

@@ 3.2 Baseline Clustering Effects
$\mathrm{logit}(p_{ijk})=\text{fixed effects}+u_{i}+v_{ij},$
@@
where $i$ indexes communities, $j$ indexes families, and $k$ indexes individual births. Here, $p_{ijk}$ denotes the conditional probability of modern prenatal care; $u_{i}\sim\mathrm{N}(0,\sigma_{u}^{2})$ captures community-level heterogeneity, while $v_{ij}\sim\mathrm{N}(0,\sigma_{v}^{2})$ models variation at the family level.
This hierarchical formulation explicitly accommodates the two main sources of dependence identified in the exploratory analysis—shared effects within communities and within families—providing a natural and interpretable framework aligned with the substantive context of the study.
To further assess the adequacy and effectiveness of our model relative to the GLMM and conventional GEE, we compare their predictive performance. The GLMM is fitted using the glmmTMB package (McGillycuddy et al.,, 2025) in R, with predicted probabilities $\hat{p}_{ijk}$ obtained via the predict function. The GEE with an exchangeable working correlation is fitted using the gee package (Carey,, 2024).
We implement a 5-fold cross-validation procedure repeated 15 times and compute two commonly used metrics: the Brier Score (Brier,, 1950) and the Log Loss, defined as
Brier Score $=\frac{1}{N}\sum_{i=1}^{n}\sum_{j=1}^{m_{i}}(y_{ij}-\hat{p}_{ij})^{2},$
where $n$ is the number of communities, $m_{i}$ the size of community $i$, NN the total number of observations, $y_{ij}$ the observed outcome, and $\hat{p}_{ij}$ the predicted probability.
For each repetition, the performance is averaged over the 5 folds:
and the overall cross-validated score is computed as
$\text{CV}_{l}=\frac{1}{15}\sum_{r=1}^{15}\bar{s}_{r,l},\quad l=1,2,$
where $r$ indexes repetitions and $k$ indexes folds.
@@
To assess relative performance, we perform pairwise $t$-tests comparing the mean CV score differences between our model (11) and the alternative models. The corresponding $t$-statistics and $p$-values are reported in Table 2.
Estimate $\textbf{CV}_{1}$ $\textbf{CV}_{2}$
GEE (Exch) GLMM GEE (Exch) GLMM
$t$-Statistics -9.9055 -51.195 -11.685 -49.645
pp-Value 0.0000 0.0000 0.0000 0.0000
Table 2: Model comparison for modern prenatal care data.
@@
The results indicate that our model (11) attains the lowest mean CV scores with minimal variability across both evaluation metrics, demonstrating superior predictive performance compared with the alternative approaches. In contrast, the GLMM exhibits notably higher mean CV scores and greater variability, reflecting lower predictive accuracy and increased sensitivity to data perturbations. As reported in Table 2, all $p$-values are below 0.0001, providing strong statistical evidence that the predictive advantage of our model over the competitors is highly significant.
3.3 Refined Analysis of the Correlation Models
Beyond community- and family-level clustering, other features may induce additional correlations. For example, the ethnic-linguistic identifier indig links multiple birth records, serving as a natural ethnic-linguistic grouping variable.
To illustrate the flexibility and practical utility of our proposed approach, we extend the baseline correlation model (11) to incorporate socio-demographic and household-level covariates relevant to prenatal care choices:
$+\alpha_{12}\,w_{ijk,12}+\alpha_{13}\,w_{ijk,13},$ (12)
where the indices $i,j,k$ follow the same convention as in (11).
The additional covariates are defined as:
Beyond generic community- and family-level clustering, model (3.3) allows correlation to vary with shared ethnic-linguistic background, socioeconomic similarity, parental education differences, and household amenities—factors relevant to prenatal care access and choices.
$+\alpha_{12}\,w_{ijk,12}+\alpha_{13}\,w_{ijk,13},$ (13)
where the notation follows that of (3.3).
The parameter estimates and corresponding $p$-values for models (3.3) and (3.3) are summarized in Table LABEL:table:corr_family, facilitating comparison of the influence of generic versus covariate-specific correlation components.
Variable Model (3.3) Model (3.3)
@@ 3.3 Refined Analysis of the Correlation Models
Intercept ($\alpha_{0}$) $0.0471$. $0.0937$ 0.07390.0739* $0.0164$
mom ($\alpha_{1}$) 0.82660.8266*** $0.0000$
Ladino ($\alpha_{2}$) 0.0187-0.0187 $0.5628$ $0.0203$ $0.4893$
Unskilled ($\alpha_{5}$) $0.3592$ $0.3001$ 0.84540.8454*** $0.0000$
Professional ($\alpha_{6}$) $0.0756$ $0.7874$ 0.29300.2930** $0.0013$
Agri-self ($\alpha_{7}$) $0.0331$. $0.0823$ 0.04380.0438* $0.0102$
Agri-empl ($\alpha_{8}$) $0.0115$ $0.6601$ 0.05410.0541*** $0.0006$
Skilled ($\alpha_{9}$) $0.0629$ $0.3018$ 0.18420.1842*** $0.0000$
momEd ($\alpha_{10}$) $0.0066$ $0.6704$ 0.0279-0.0279* $0.0413$
husEd ($\alpha_{11}$) $0.0041$ $0.7744$ 0.0352-0.0352** $0.0023$
toilet ($\alpha_{12}$) $0.0073$ $0.8331$ 0.0413-0.0413 $0.2276$
TV ($\alpha_{13}$) 0.0004-0.0004 $0.9809$ 0.0186-0.0186 $0.2135$
In Model (3.3), the hierarchical clustering effects $\hat{\alpha}_{0}$ and $\hat{\alpha}_{1}$ remain significant, while most additional family-level covariates are insignificant with near-zero effects, likely because the strong generic familial clustering $w_{ijk,1}$ captures most of their potential variation.
Two exceptions are $\hat{\alpha}_{4}$ (Indigenous Spanish ethnic-linguistic group) and $\hat{\alpha}_{7}$ (husband agri-self employment), both marginally significant (p <0.1<0.1), indicating additional dependence beyond generic family- and community-level effects. Even after accounting for general family and community clustering, births in the Indigenous Spanish population are somewhat less correlated, whereas mothers in agri-self households show greater similarity in prenatal care choices. These patterns may reflect cultural norms or occupation-related constraints, suggesting potential targets for health interventions or further investigation.
For Model (3.3), which omits the generic familial clustering term $w_{ijk,1}$, several notable patterns emerge:
  1. 1.
    Effect of removing $w_{ijk,1}$: Without the generic familial clustering term, parameters for the family-level covariates generally have larger magnitudes and lower p-values compared to Model (3.3). This confirms that the model effectively partitions overall familial clustering into more detailed subcomponents.
  2. 2.
  3. 3.
  4. 4.
  5. 5.
These results reveal insights often obscured by conventional models that capture only generic clustering. While community- and family-level dependence affects prenatal care choices, the underlying social and structural drivers typically remain hidden. Model (3.3) uncovers nuanced correlation patterns linked to social, ethnic, socio-demographic, and economic factors, highlighting mechanisms behind observed clustering and suggesting targets for health interventions or further investigation.
To examine individual-level factors that may influence correlations in prenatal care choices more closely, we extend Model (3.3) by incorporating pairwise differences in selected individual covariates:
where $i,j,k$ follow the same indexing as in (11). The additional covariates are defined as
$w_{ijk,16}=|\mathrm{birthOrd}_{ij}-\mathrm{birthOrd}_{ik}|,$
representing pairwise differences in maternal age, child age, and birth order, respectively. All other covariates remain as specified in (11) and (3.3). The corresponding parameter estimates and p-values are summarized in Table LABEL:table:corr_full.
Variable Estimate P-value
Intercept ($\alpha_{0}$) 0.06830.0683* $0.0283$
mom ($\alpha_{1}$) 0.82690.8269*** $0.0000$
Ladino ($\alpha_{2}$) 0.0172-0.0172 $0.6092$
Indigenous non-Spanish ($\alpha_{3}$) 0.0034-0.0034 $0.9389$
Indigenous Spanish ($\alpha_{4}$) 0.0648-0.0648. $0.0792$
Unskilled ($\alpha_{5}$) $0.3768$ $0.2653$
Professional ($\alpha_{6}$) $0.0944$ $0.7185$
Agri-self ($\alpha_{7}$) $0.0293$ $0.1467$
Agri-empl ($\alpha_{8}$) $0.0166$ $0.5786$
Skilled ($\alpha_{9}$) $0.0578$ $0.3209$
momEd ($\alpha_{10}$) $0.0095$ $0.5445$
husEd ($\alpha_{11}$) $0.0027$ $0.8541$
toilet ($\alpha_{12}$) $0.0061$ $0.8722$
TV ($\alpha_{13}$) $0.0021$ $0.8981$
motherAge ($\alpha_{14}$) $0.0072$ $0.6238$
childAge ($\alpha_{15}$) 0.0251-0.0251 $0.1906$
birthOrd ($\alpha_{16}$) 0.0130-0.0130 $0.3412$
Analysis of Model (3.3) yields results broadly consistent with our earlier findings. Although none of the additional individual-level covariates reach statistical significance, both $\hat{\alpha}_{15}$ and $\hat{\alpha}_{16}$ are estimated to be negative. This suggests a gradual weakening of shared prenatal care patterns over time and reduced similarity in care choices among children with larger age differences in larger families. Such patterns may reflect temporal changes in access, preferences, or health policies, as well as generational differences within multi-child households. However, given the limited data for these specific covariates, these observations should be interpreted cautiously, and further investigation is warranted before drawing definitive conclusions.
3.4 Comparisons with GLMM
We assess the performance of the GLMM as a main competing approach in this context. In addition to the community- and family-level effects described in Section 3.2, we explore extensions incorporating additional hierarchical random effects. The GLMM is initially fitted using the lme4 package (Bates et al.,, 2015). For more complex specifications, where convergence issues arise due to model complexity, we employ the glmmTMB package, which is generally more robust for large models with multiple random effects. The corresponding results are presented in Tables 5 and 6.
Estimate Community Family indig husEmpl momEd husEd
Std.Dev. $0.0084$ $24.2176$ $0.0119$ $1.0357$ $0.0046$ $0.0020$
$p$-Value NA NA $0.4950$ NA NA NA
Table 5: Estimates of the random effects by the GLMM.
Estimate toilet TV motherAge childAge birthOrd
Std.Dev. $0.0127$ $0.0019$ $1.0726$ $3.3576$ $3.8209$
$p$-Value NA $1.0000$ $0.8731$ NA $1.0000$
Table 6: Estimates of the random effects by the GLMM.
3.5 Additional results
@@
We further illustrate the broad applicability of our framework with two additional real data analyses—a longitudinal binary dataset and a clustered count dataset—together with extensive simulation studies. Across these settings, our method consistently outperforms competitors such as the GLMMs in both predictive accuracy and interpretability, demonstrating clear competitive advantages in practical applications. Complete details are provided in the Supplementary Material.
@@ 4 Discussion
5 Appendices
All technical details, as well as two additional real data analyses—including a longitudinal binary dataset from a toenail infection clinical trial and a clustered count health dataset—are provided in the Supplementary Material. Extensive simulation results illustrating the performance of the proposed method under various scenarios are also included therein.
@@ 6 Data Availability Statement
Appendix A Theory
Regularity conditions. We assume:
  • (A1)
  • (A2)
  • (A3)
Theorem 1 (Theorem 4, Xie and Yang, (2003)).
where ${\bf M}_{1,n}={\bf H}_{1}(\boldsymbol{\alpha}_{0},\boldsymbol{\beta}_{0},\phi_{0})$ is defined in Section 2.3 of the main part.
  • (B1)
  • (B2)
  • (B3)
Theorem 2.
The exact expression of the leading matrix component of ${\bf M}_{2,n}$ matches ${\bf H}_{2}$ defined in Section 2.3 of the main part. However, due to the high complexity of the Hessian matrix, we employ the numerical method to obtain its approximation. Denote
where ${\bf S}_{2}$ is the score function defined in Section 2.3 of the main part and $h\in\mathbb{R}^{+}$ denotes a sufficiently small constant. Let ${\bf A}(h)=({\mathbf{a}}_{1}(h),{\mathbf{a}}_{2}(h),\dots,{\mathbf{a}}_{d}(h))$. Lemma 4.2.2 in Dennis Jr and Schnabel, (1996) shows that, under mild conditions,
Discussion. Theorems 1 and 2 together show that the proposed two-stage procedure enjoys desirable large-sample guarantees. In particular, the estimation of $\boldsymbol{\beta}$ is asymptotically insensitive to small misspecification of $\boldsymbol{\alpha}$, while the pseudo-likelihood estimator of $\boldsymbol{\alpha}$ remains consistent due to the identifiability of the Gaussian working model. Importantly, these results hold under bounded cluster sizes and fixed covariate dimensions; extensions to high-dimensional or diverging-cluster settings would require additional technical tools, which we leave for future work.
Appendix B Proof of Theorem 2
Proof of Theorem 2.
By the strong law of large numbers, we have
Denote
[…]
By our assumptions and the central limit theorem, we have
[…]
and thus we have
by Slutsky’s theorem and the symmetry of the Gaussian distribution. ∎
Appendix C Calculation of the Pseudo-expectation
We calculate the pseudo-expectation $\tilde{\mathbb{E}}({\bf J}_{i})$ defined in Section 2.3 of the main part. We need the following lemma:
@@ Appendix C Calculation of the Pseudo-expectation
Appendix D Simulations

### v2

@@ 3.3 Comparisons with GLMM
$\mathrm{logit}(p_{ijklt})=\text{fixed effects}+u_{i}+v_{ij}+s_{ik}+c_{il},$ (15)
@@
where $i$ indexes communities, $j$ indexes families, $k$ indexes ethnic-linguistic identities, $l$ indexes husband’s employment status, and $t$ indexes individual births. Here, $p_{ijklt}$ denotes the conditional probability of modern prenatal care; $u_{i}\sim\mathrm{N}(0,\sigma_{u}^{2})$ captures community-level heterogeneity; $v_{ij}\sim\mathrm{N}(0,\sigma_{v}^{2})$ captures family-level heterogeneity; $s_{ik}\sim\mathrm{N}(0,\sigma_{s}^{2})$ models variation of the ethnic-linguistic identity; $c_{il}\sim\mathrm{N}(0,\sigma_{c}^{2})$ models variation of the husband’s employment status. This hierarchical formulation explicitly accommodates the two main sources of dependence and the additional covariate-based correlations, providing a natural and interpretable framework aligned with the substantive context of the study.
However, fitting the GLMM (15) yields statistically insignificant estimates for $\sigma_{s}$ and $\sigma_{c}$, which goes against our former empirical findings and analyzing results, suggesting that the model could not discern the variability attributable to ethnic-linguistic identity or husband’s employment status. This apparent discrepancy with our initial empirical findings arises because the GLMM can only estimate a single variance component across all occupations (as a random effect), while our method directly parameterizes the correlation structure itself. This allows us to estimate distinct coefficients (e.g., $\alpha_{5}$ for agri-self in model (14)) that quantify how each specific occupation category independently modulates the strength of within-cluster dependence. Consequently, our proposed method enhances interpretability of how specific covariates influence the correlation, and enables fine-grained analysis, which is often infeasible with conventional GEE and GLMM methods.
To substantiate the claim that including indig and husEmpl is necessary for capturing the correlation structure, we next assess the improvement in model fit. To empirically validate this improvement and rigorously benchmark our approach, we compare its predictive performance against our initial model (12), the GLMM and GEE alternatives. A reduced GLMM is also fitted for comparison with our initial model (12):
$\mathrm{logit}(p_{ijt})=\text{fixed effects}+u_{i}+v_{ij},$ (16)
where the notation follows that of (15). Predicted probabilities for the GLMMs are obtained using the predict function. Additionally, we fit a GEE with an exchangeable working correlation structure. Both the GLMM and GEE models incorporate the same set of fixed effects as specified in (3.1).
@@
We implement a 5-fold cross-validation procedure repeated 15 times, and the model performance is evaluated using the Brier Score (Brier,, 1950) and the Log Loss, with $\text{CV}_{1}$ and $\text{CV}_{2}$ denoting the mean cross-validated scores for the Brier Score and Log Loss, respectively; see Supplementary Material for their definitions. For both metrics, smaller values indicate better predictive performance. To assess relative performance, we perform pairwise $t$-tests comparing the mean CV score differences between our model (14) and the alternative models. The corresponding $t$-statistics and $p$-values are reported in Table 6.
Criteria Model (12) GEE (Exch) GLMM (15) GLMM (16)
$\textbf{CV}_{1}$ 0.0001(-5.2233) 0.0000(-9.2625) 0.0000(-54.614) 0.0000(-50.270)
$\textbf{CV}_{2}$ 0.0010(-4.1294) 0.0000(-9.6811) 0.0000(-52.560) 0.0000(-48.618)
Table 6: Model comparison for toenail infection data.
@@
The results demonstrate that our proposed model (14) achieves the lowest mean cross-validation scores with minimal variance across both evaluation metrics, indicating superior predictive performance over all alternative approaches. In contrast, the GLMM yields substantially higher mean error scores accompanied by greater variability, suggesting not only reduced predictive accuracy but also higher sensitivity to variations in the data. As summarized in Table 6, all pairwise comparisons yield $p$-values below 0.01, providing strong statistical evidence that the performance advantage of our model is highly significant.
Furthermore, while our initial baseline model (12) has already outperformed conventional methods, incorporating the covariates indig and husEmpl in model (14) leads to a statistically significant improvement in predictive performance, underscoring the value of explicitly modeling covariate-dependent correlation structures.
4 Additional Real Data Analysis
Variable Description
patientID Unique identifier for each patient in the trial
time The time in month when the visit actually took place: $[0,18.5]$
visit Number of visit attended: $\{1,2,3,4,5,6,7\}$
Table 7: Description of variables of the toenail infection dataset.
Parameter estimates for both models, along with their corresponding $p$-values, are reported in Table 8.
Variable Model (18) Model (19)
@@ 4 Additional Real Data Analysis
treatment ($\beta_{1}$) $0.0313$ $0.8900$ $0.0292$ $0.9014$
time ($\beta_{2}$) 0.1530-0.1530*** $0.0000$ 0.1703-0.1703*** $0.0000$
Notes: . for $\text{p}<0.1$, * for $\text{p}<0.05$, ** for $\text{p}<0.01$ and *** for $\text{p}<0.001$.
Table 8: Parameter estimates and corresponding p-values for (18) and (19).
For the correlation models, the results confirm a positive correlation between repeated measurements from the same patient, as captured by $\alpha_{0}$ or $\tilde{\alpha}_{0}$. The small estimates of $\alpha_{1}$ and $\tilde{\alpha}_{1}$ indicate that receiving the same treatment does not introduce additional correlation beyond the time effect, suggesting that the primary source of correlation is the patient-specific effect—an intuitively reasonable finding in this clinical context. This correlation decreases as the time gap between measurements increases, as reflected by the negative estimates of $\alpha_{2}$ or $\tilde{\alpha}_{2}$.
In this relatively standard setting, conventional methods such as the GEE and GLMM are directly applicable. For the GEE, we consider two working correlation structures: exchangeable and AR(1). To ensure stable estimation of the AR(1)-GEE model in gee, patients with only a single recorded visit are excluded from the analysis.
$\mathrm{logit}(p_{ijk})=\text{fixed effects}+u_{i}+v_{j}+w_{k},$
where $p_{ijk}=\mathbb{E}(y_{ijk})$ is the expected response for the $i$th patient, $j$th treatment type, and $k$th visit. Here, $u_{i}\sim\mathrm{N}(0,\sigma_{u}^{2})$ captures patient-level heterogeneity, $v_{j}\sim\mathrm{N}(0,\sigma_{v}^{2})$ captures treatment-level variation, and $w_{k}\sim\mathrm{N}(0,\sigma_{w}^{2})$ models visit-specific effects. Unlike our proposed method, the standard GLMM framework lacks a mechanism to model correlations directly based on continuous variables, unless they are included to have random slopes, which drastically changes the question to be addressed. Thus, to account for temporal correlation, it is restricted to using the discrete variable visit.
Fitting the models with both alternative approaches yields quantitatively similar results. We then compare their predictive performance against our proposed method. To assess predictive accuracy, we employ a 15-repeat stratified 5-fold cross-validation scheme. The dataset is first partitioned into two mutually exclusive subsets by treatment type, with random sampling conducted within each subset to preserve the original treatment distribution. The same cross-validation performance evaluating metrics is used as in Section 3.
Pairwise $t$-tests comparing mean CV score differences between model (19) and each of the other four models yield the $p$-values and $t$-statistics reported in Table 9. With the sole exception of the $\text{CV}_{1}$ comparison against the GEE (Exch) model, all differences are statistically significant negative at the 0.05 level. From an applied perspective, these results mean that our approach not only predicts patient outcomes more accurately but also clarifies how the timing of repeated measurements shapes correlation patterns over the treatment period. This combination of predictive strength and interpretability provides actionable insights for clinical trial design and patient monitoring strategies, offering a clear advantage over competing methods in both practical utility and statistical performance—likely attributable to its superior adequacy in model fitting.
Criteria Model (18) GEE (Exch) GEE (AR-1) GLMM
$\textbf{CV}_{1}$ 0.0004(-4.6233) 0.7984(-0.2603) 0.0003(-4.7688) 0.0000(-64.382)
$\textbf{CV}_{2}$ 0.0011(-4.1041) 0.0346(-2.3408) 0.0000(-6.5205) 0.0000(-27.065)
Table 9: Model comparison for toenail infection data.
@@
We further illustrate the broad applicability of our framework with additional analysis of the motivating modern prenatal care dataset and analysis of a clustered health count dataset, together with extensive simulation studies. Across these settings, our method consistently outperforms competitors such as the GLMMs in both predictive accuracy and interpretability, demonstrating clear competitive advantages in practical applications. Complete details are provided in the Supplementary Material.
@@ 6 Data Availability Statement
7 Appendices
All technical details, as well as the additional real data analyses, are provided in the Supplementary Material. Extensive simulation results illustrating the performance of the proposed method under various scenarios are also included therein.
@@ 7 Appendices
Appendix A Algorithm
A.1 Modified Fisher’s Scoring Algorithm
To jointly estimate $(\boldsymbol{\alpha},\boldsymbol{\beta},\phi)$, we employ an iterative procedure based on a modified Fisher scoring algorithm that alternates between updates of $\boldsymbol{\beta}$, $\boldsymbol{\alpha}$, and $\phi$, combining the score functions from the GEE with the pseudo-likelihood. Define the score functions as
${\bf A}=({\bf Q}\otimes{\bf Q})\boldsymbol{\Xi}({\bf Q}\otimes{\bf Q})^{\top},$
where $\boldsymbol{\Xi}$ is diagonal with entries
Algorithm 1 Modified Fisher’s Scoring Algorithm
0: Estimates $(\boldsymbol{\alpha},\boldsymbol{\beta},\phi)$
1:repeat
2:  Update $\phi^{(k)}$ using $\boldsymbol{\beta}^{(k-1)}$
4:  repeat
5:   Update
6:   Increment $s\leftarrow s+1$
7:  until convergence; update $\boldsymbol{\alpha}^{(k)}$
8:  Update
9:  Increment $k\leftarrow k+1$
10:until convergence
High-dimensional $\boldsymbol{\eta}_{i}$ can slow convergence. To mitigate this, we replace ${\bf J}_{i}$ with a pseudo-expectation $\tilde{\operatorname{\mathbb{E}}}({\bf J}_{i})$ capturing the dominant first-order structure while ignoring higher-order dependencies (see Section A.2), yielding stable estimates with negligible bias at lower cost.
A.2 Calculation of the Pseudo-expectation
We calculate the pseudo-expectation $\tilde{\mathbb{E}}({\bf J}_{i})$ defined in Section A.1. We need the following lemma:
@@ A.2 Calculation of the Pseudo-expectation
Appendix B Theory
B.1 Asymptotic Property
Regularity conditions. We assume:
  • (A1)
  • (A2)
  • (A3)
Theorem 2 (Theorem 4, Xie and Yang, (2003)).
where ${\bf M}_{1,n}={\bf H}_{1}(\boldsymbol{\alpha}_{0},\boldsymbol{\beta}_{0},\phi_{0})$ is defined in Section A.1.
  • (B1)
  • (B2)
  • (B3)
Theorem 3.
The exact expression of the leading matrix component of ${\bf M}_{2,n}$ matches ${\bf H}_{2}$ defined in Section A.1. However, due to the high complexity of the Hessian matrix, we employ the numerical method to obtain its approximation. Denote
where ${\bf S}_{2}$ is the score function defined in Section A.1 and $h\in\mathbb{R}^{+}$ denotes a sufficiently small constant. Let ${\bf A}(h)=({\mathbf{a}}_{1}(h),{\mathbf{a}}_{2}(h),\dots,{\mathbf{a}}_{d}(h))$. Lemma 4.2.2 in Dennis Jr and Schnabel, (1996) shows that, under mild conditions,
Discussion. Theorems 2 and 3 together show that the proposed two-stage procedure enjoys desirable large-sample guarantees. In particular, the estimation of $\boldsymbol{\beta}$ is asymptotically insensitive to small misspecification of $\boldsymbol{\alpha}$, while the pseudo-likelihood estimator of $\boldsymbol{\alpha}$ remains consistent due to the identifiability of the Gaussian working model. Importantly, these results hold under bounded cluster sizes and fixed covariate dimensions; extensions to high-dimensional or diverging-cluster settings would require additional technical tools, which we leave for future work.
B.2 Proof of Theorem 3
Proof of Theorem 3.
By the strong law of large numbers, we have
Denote
[…]
By our assumptions and the central limit theorem, we have
[…]
and thus we have
by Slutsky’s theorem and the symmetry of the Gaussian distribution. ∎
Appendix C Simulations